### v3

# Universal finite-size scaling in high-dimensional critical phenomena

Abstract
We present a new unified theory of critical finite-size scaling for lattice statistical mechanical models with periodic boundary conditions above the upper critical dimension. Our theory is based on recent mathematically rigorous results for linear and branched polymers, multi-component spin systems, and percolation. Both short-range and long-range interactions are included. The universal finite-size scaling is inherited from the scaling of the system unwrapped to the infinite lattice. We also present conjectures for universal scaling profiles for the susceptibility and two-point function plateau in a critical window. For free boundary conditions, the universal scaling has been proven to apply at a pseudocritical point for hierarchical spins, and we conjecture that this holds generally.
I Introduction
The subject of critical phenomena for lattice models in statistical mechanics is a cornerstone of theoretical physics. The critical behaviour of the Ising and multi-component $|\varphi|^{4}$ models for ferromagnetism, of percolation, and of linear and branched polymers, forms a central part of the theory. The rich and universal fractal geometry connected with second-order phase transitions in these models, and its characterisation in terms of critical exponents, is an ongoing source of fascination both in physics and in mathematics.
It has long been understood that the dependence on the spatial dimension is an important feature in critical phenomena. In particular, there is typically an upper critical dimension $d_{\mathrm{c}}$ above which mean-field critical exponents occur. For models with short-range (SR) interactions, $d_{\mathrm{c}}=4$ for Ising and $|\varphi|^{4}$ spin systems, $d_{\mathrm{c}}=4$ for linear polymers (self-avoiding walk), $d_{\mathrm{c}}=6$ for percolation, and $d_{\mathrm{c}}=8$ for branched polymers (lattice trees and animals).
Since laboratory samples and simulation experiments involve finite systems, the finite-size scaling (FSS) of phase transitions forms an important part of the theory. Above the upper critical dimension, the role played by boundary conditions in FSS has been a subject of some controversy and much discussion, e.g., [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13]. Some of the history is recounted in [9]. It is therefore useful to have mathematically rigorous results for FSS.
In this work, we present a new and unified theory of FSS under periodic boundary conditions (PBC), based on unwrapping the finite model to the infinite lattice as in Figure 1. A general comparison principle between the original and the unwrapped systems is proposed and verified for the Ising model, the self-avoiding walk, percolation, and branched polymers above their upper critical dimensions, using the rigorous results of [14, 15, 16, 17, 18]. Thereby, we obtain mathematically rigorous results on the finite-size critical exponents for these models under PBC.
Figure 1: The unwrapping of a lattice tree from the 2-dimensional discrete torus of period 5 to the infinite lattice.
For all of the SR models mentioned above, we review recent proofs that in a volume $V$ with PBC in dimensions $d>d_{\mathrm{c}}$, and at the infinite-volume critical point, the susceptibility has size (at least) $V^{\frac{2}{d_{\mathrm{c}}}}$, and the two-point function has a plateau of size (at least) $V^{\frac{2}{d_{\mathrm{c}}}-1}$. We also identify a critical window of width (or rounding scale) $V^{-\frac{2}{\gamma d_{\mathrm{c}}}}$ around the infinite-volume critical point, where $\gamma$ is the critical exponent for the infinite-volume susceptibility.
Long-range (LR) versions of the models have couplings decaying like $r^{-(d+\alpha)}$, with $\alpha\in(0,2)$ for $d\geq 2$ and $\alpha\in(0,1)$ for $d=1$ (the enhanced restriction for $d=1$ is to ensure a second-order phase transition). LR models have upper critical dimensions $d_{\mathrm{c},\alpha}=\frac{\alpha}{2}d_{\mathrm{c}}$, with $d_{\mathrm{c}}$ the SR upper critical dimension. For LR models, $d>d_{\mathrm{c},\alpha}$ includes low dimensions, even $d=1$, when $\alpha$ is small enough. Our general theory also applies to LR models above $d_{\mathrm{c},\alpha}$. It is always the short-range $d_{\mathrm{c}}$, not the long-range $d_{\mathrm{c},\alpha}$, that appears in the finite-size scaling exponents mentioned in the previous paragraph. We demonstrate this with the long-range self-avoiding walk, using a recent result of [19].
[…]
Throughout the paper, we discuss short-range and long-range models simultaneously. Formulas involving $\alpha$ apply to LR models for any choice of $\alpha\in(0,\min\{d,2\})$, and apply to SR models after setting $\alpha=2$.
Notation. For dimension $d\geq 1$, the hypercubic lattice is denoted $\mathbb{Z}^{d}$, and the discrete torus of period $R$ and volume $V=R^{d}$ is denoted $\mathbb{T}_{R}^{d}$. We sometimes identify a point $x$ in the torus with its representative in $\frac{R}{2},\frac{R}{2})^{d}\cap\mathbb{Z}^{d}$.
We write $f\lesssim g$ or $f=O(g)$ to mean there is a constant $C>0$ such that $f\leq Cg$, write $f\ll g$ to mean $\lim f/g=0$, and write $f\asymp g$ to mean $f\lesssim g\lesssim f$. We write $|x|=\max\{|x_{1}|,\ldots,|x_{d}|\}$ for the $\ell^{\infty}$ norm, and $|x|\!|\!|=\max\{|x|,1\}$ to avoid division by 0.
Critical exponents. For a given model, let $\beta_{\mathrm{c}}$} denote the infinite-volume critical point, and let $t=({\beta_{\mathrm{c}}}-\beta)/{\beta_{\mathrm{c}}}>0$ denote the reduced inverse temperature. The two-point functions on $\mathbb{Z}^{d}$ and $\mathbb{T}_{R}^{d}$ are denoted, respectively, by $G_{\beta}(x)$ and $G_{\beta,R}(x)$. Let $\chi(\beta)=\sum_{x\in\mathbb{Z}^{d}}G_{\beta}(x)$ denote the susceptibility on $\mathbb{Z}^{d}$}, and let $\chi_{R}(\beta)=\sum_{x\in\mathbb{T}_{R}^{d}}G_{R,\beta}(x)$ denote the susceptibility on $\mathbb{T}_{R}^{d}$. The two-point functions and susceptibilities are not truncated. We assume that both two-point functions are non-decreasing functions of $\beta$. Let $\xi(\beta)$ denote the correlation length on $\mathbb{Z}^{d}$}. We do not use a correlation length on the torus. The $\mathbb{Z}^{d}$} critical exponents $\eta,\gamma,\nu$ are defined by
$G_{{\beta_{\mathrm{c}}}}(x)\asymp\frac{1}{|x|^{d-2+\eta}}\qquad(|x|\to\infty),$ (1)
$\chi(\beta)\asymp t^{-\gamma},\qquad\xi(\beta)\asymp t^{-\nu}\qquad(t\downarrow 0).$ (2)
Above the upper critical dimension $d_{\mathrm{c},\alpha}$, the critical exponents are proved in many cases (e.g., [20, 21, 22, 23, 24, 25, 26, 27]) to take their mean-field values $\eta=2-\alpha$, $\nu=\gamma/\alpha$, and
$\gamma=\begin{cases}1&(\text{self-avoiding walk, spin, percolation})\\ \frac{1}{2}&(\text{lattice trees and animals}).\end{cases}$ (3)
Our main rigorous result (Theorem 1 in Section II) gives a sufficient condition for the short- or long-range torus susceptibility and two-point function to obey
[…]
uniformly in large $R$, with the short-range $d_{\mathrm{c}}$. The constant “plateau” term in the lower bound of $G_{R,{\beta_{\mathrm{c}}}}(x)$ dominates $|x|^{-(d-\alpha)}$ as soon as
$|x|\gg R^{p}\;\;\text{with}\;\;p=\frac{d-\alpha\frac{d}{d_{\mathrm{c},\alpha}}}{d-\alpha}.$ (5)
Note that the exponent $p$ is less than $1$ for $d>d_{\mathrm{c},\alpha}$, so the plateau dominates over all but a vanishingly small proportion of the torus. The plateau term contributes $V^{\frac{2}{d_{\mathrm{c}}}}=R^{\frac{2d}{d_{\mathrm{c}}}}=R^{\frac{\alpha d}{d_{\mathrm{c},\alpha}}}$ to the susceptibility, which is much larger than the contribution $R^{\alpha}$ from the decaying term in the two-point function. Our proof also identifies the width of the critical window to be $V^{-\frac{2}{\gamma d_{\mathrm{c}}}}$, and it gives matching upper bounds for (4) at the edge (high-temperature, disordered side) of the critical window. Theorem 1 is proved in Section III, and in Section IV we verify its sufficient condition for the case of LR self-avoiding walk. Further applications of Theorem 1 are discussed in Section V.
Beyond the computation of FSS critical exponents, more precise information about the scaling of the susceptibility and two-point function plateau can be sought. For the $n$-component $|\varphi|^{4}$ model in dimensions $d>4$ with PBC, an exact profile for the amplitude of the susceptibility in the window was computed in 1985 using a Wilsonian renormalisation group method [28, 29]; we review this computation in Appendix D. A mathematically rigorous derivation of the same profile has been carried out recently on the hierarchical lattice [30, 31] for dimensions $d\geq 4$ (including $d=d_{\mathrm{c}}=4$). For $n=1$, the profile is the same as its analogue on the complete graph (Curie–Weiss model). Despite the results of [28], the profile has subsequently received scant attention in the FSS literature, for example it is not explicitly mentioned in any of [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13]. Part of our purpose is to focus attention on the importance of the profile. The profiles for self-avoiding walk and branched polymers on the complete graph have been computed, and we conjecture that these profiles also apply to SR and LR models at and above their critical dimensions. We also discuss a profile for percolation, which is of a different character than the others. Our conjectures for the universal profiles are stated in Section VI.
For the hierarchical $n$-component $|\varphi|^{4}$ model with $d\geq d_{\mathrm{c}}=4$, it has been proved that, under FBC in volume $V=R^{d}$, the universal behaviour observed for PBC in the critical window about the infinite-volume critical point is exactly duplicated around a pseudocritical point which is shifted from the infinite-volume critical point by $R^{-1/\nu}=R^{-2}$ for $d>4$ (this shift is also observed, e.g., in [32, 4]) and by $R^{-2}(\log R)^{\frac{n+2}{n+8}}$ for $d=4$ [30, 31]. The critical windows for PBC and FBC do not overlap when $d\geq d_{\mathrm{c}}$. In Section VII, we formulate a conjecture that similar behaviour holds more generally.
II General plateau theorem
Our general plateau theorem holds under two hypotheses. The hypotheses have been verified for many examples, discussed in Sections IV and V. Our first hypothesis concerns the $\mathbb{Z}^{d}$} two-point function $G_{\beta}(x)$ near criticality.
Hypothesis 1.
The susceptibility obeys (2) with critical exponent $\gamma>0$. There is a correlation length satisfying $\xi(\beta)\asymp t^{-\nu}$ with $\nu=\gamma/\alpha$, a function $g:[0,\infty)\to[0,\infty)$ with $g(u)\lesssim(1+u)^{-(\alpha+\varepsilon)}$ for some $\varepsilon>0$, and a constant $s_{1}>0$ such that
[…]
Hypothesis 1 asserts that $G_{\beta}(x)$ has its critical decay $|x|^{-(d-\alpha)}$ up to the scale of the correlation length, beyond which it decays more rapidly; this is essentially a definition of a correlation length. In applications, the function $g$ takes the form
$g_{\mathrm{SR}}(u)=e^{-cu},\qquad g_{\mathrm{LR}}(u)=\frac{1}{(1+cu)^{2\alpha}}$ (8)
for SR and LR models respectively.
Our second hypothesis concerns the torus two-point function $G_{R,\beta}(x)$. For its statement, we introduce the unwrapped two-point function
$\Gamma_{R,\beta}(x)=\sum_{u\in\mathbb{Z}^{d}}G_{\beta}(x+Ru)\qquad(x\in\mathbb{T}_{R}^{d}),$ (9)
which can be considered either as a function on the torus or as a periodic function on $\mathbb{Z}^{d}$. (Recall we identify $\mathbb{T}_{R}^{d}$ with $\frac{R}{2},\frac{R}{2})^{d}\cap\mathbb{Z}^{d}$.) The sum is over all possible unwrapped locations $x+Ru$ of the torus point $x\in\mathbb{T}_{R}^{d}$. The unwrapped two-point function is inspired by the unwrapping shown in Figure 1.
Hypothesis 2.
There are constants $c_{0},s_{2}>0$ such that
$\Big(1-c_{0}\frac{\chi(\beta)^{d_{\mathrm{c}}/2}}{V}\Big)\Gamma_{R,\beta}(x)\leq G_{R,\beta}(x)\leq\Gamma_{R,\beta}(x),$ (10)
uniformly in $x\in\mathbb{T}_{R}^{d}$ and in $\beta,R$ satisfying $\xi(\beta)\geq s_{2}R$.
Hypothesis 2 is a comparison principle between the torus and unwrapped models. The upper bound of (10) reflects lesser interaction in the unwrapped model, and in applications it holds in any dimension. The lower bound of (10) quantifies the error in the comparison, and it is expected to hold only above the upper critical dimension $d_{\mathrm{c},\alpha}$. We emphasise it is $\chi^{d_{\mathrm{c}}/2}$, not $\chi^{d_{\mathrm{c},\alpha}/2}$, that occurs in (10), even for LR models. The power $d_{\mathrm{c}}/2$ arises from the topology of certain Feynman diagrams, which are identical for SR and LR models (see Appendix B).
Theorem 1.
Let $d>d_{\mathrm{c},\alpha}$ and let $R$ be sufficiently large. Under Hypotheses 1 and 2, there is a constant $c_{1}>0$ such that, at $\beta_{*}$ defined by $t_{*}=({\beta_{\mathrm{c}}}-\beta_{*})/{\beta_{\mathrm{c}}}=c_{1}V^{-\frac{2}{\gamma d_{\mathrm{c}}}}$,
[…]
uniformly in $R$ and in $x\in\mathbb{T}_{R}^{d}$.
By the monotonicity in $\beta$, Theorem 1 immediately implies the lower bounds at $\beta_{\mathrm{c}}$} claimed in (4). Theorem 1 identifies the exponents mentioned at the end of Section I: the powers $V^{-\frac{2}{\gamma d_{\mathrm{c}}}},V^{\frac{2}{d_{\mathrm{c}}}},V^{-(1-\frac{2}{d_{\mathrm{c}}})}$ are the window scale, susceptibility scale, and plateau scale, respectively. See also Table 1.
III Proof of plateau theorem
An ingredient in the proof of Theorem 1 is the following lemma, whose elementary proof is given in Appendix A.
Lemma 2.
Let $d\geq 1$, $R\geq 1$, $a>0$ and $\xi>0$. Suppose $g:[0,\infty)\to[0,\infty)$ satisfies $g(u)\leq(1+u)^{-(a+\varepsilon)}$ for some $\varepsilon>0$. Then there is a constant $C=C(d,a,\varepsilon)>0$ such that
$\sum_{u\in\mathbb{Z}^{d}:u\neq 0}\frac{1}{|x+Ru|^{d-a}}g\bigg(\frac{|x+Ru|}{\xi}\bigg)\leq C\frac{\xi^{a}}{R^{d}}$
for all $x\in{\mathbb{Z}^{d}}$ with $|x|\leq R/2$.
The next proposition applies Lemma 2 to conclude bounds on the unwrapped two-point function $\Gamma_{R,\beta}(x)$.
Proposition 3.
Under Hypothesis 1, and assuming that $\xi(\beta)\geq\frac{3}{2s_{1}}R$ for the lower bound in (12), we have
[…]
uniformly in $x\in\mathbb{T}_{R}^{d}$.
Proof.
The upper bound is a direct consequence of Lemma 2 (with $a=\alpha$) and Hypothesis 1 ($\nu=\gamma/\alpha$ implies that $\xi^{\alpha}\asymp\chi$):
[…]
For the lower bound, we restrict the sum in (9) to $|u|\leq M$ with a well-chosen $M\geq 1$. We then wish to apply (7) to all $G_{\beta}(x+Ru)$ with $|u|\leq M$. Since $|x|\leq R/2$, by the triangle inequality we can do so if
$M\geq 1,\qquad|x+Ru|\leq\frac{R}{2}+RM\leq s_{1}\xi.$ (14)
We choose $M=\frac{2s_{1}}{3}\xi/R$ so that both inequalities assert $R\leq\frac{2s_{1}}{3}\xi$, as we have assumed to hold for the lower bound. Since $|x+Ru|\leq\frac{R}{2}+R|u|\leq\frac{3}{2}R|u|$, (7) gives
$\sum_{u\neq 0}G_{\beta}(x+Ru)$ $\gtrsim\sum_{1\leq|u|\leq M}\frac{1}{|x+Ru|^{d-\alpha}}$
$\geq\sum_{1\leq|u|\leq M}\frac{1}{(\frac{3}{2}R|u|)^{d-\alpha}}\asymp\frac{M^{\alpha}}{R^{d-\alpha}}.$ (15)
Our choice $M=\frac{2s_{1}}{3}\xi/R$, the lower bound (7) on $G_{\beta}(x)$, and $\xi^{\alpha}\asymp\chi$, then give
[…]
This completes the proof. ∎
Proof of Theorem 1.
We start by fixing $c_{1}$ large enough so that the subtracted term in the lower bound of Hypothesis 2 is harmless at $\beta_{*}$. Since $\chi(\beta)\asymp t^{-\gamma}$ by (2), we can and do fix $c_{1}$ sufficiently large so that
$c_{0}\frac{\chi(\beta_{*})^{d_{\mathrm{c}}/2}}{V}\leq c_{0}^{\prime}\frac{1}{Vt_{*}^{\gamma d_{\mathrm{c}}/2}}=\frac{c_{0}^{\prime}}{c_{1}^{\gamma d_{\mathrm{c}}/2}}\leq\frac{1}{2}.$ (17)
Concerning the restrictions $\xi(\beta_{*})\geq s_{2}R$ and $\xi(\beta_{*})\geq\frac{3}{2s_{1}}R$ required by Hypothesis 2 and Proposition 3, we use $\xi(\beta)\asymp t^{-\nu}=t^{-\gamma/\alpha}$ by (2) and use $d>d_{\mathrm{c},\alpha}$, to see that for $R$ sufficiently large we have
$\xi(\beta_{*})\asymp\frac{1}{t_{*}^{\gamma/\alpha}}=\frac{V^{\frac{2}{\alpha d_{\mathrm{c}}}}}{c_{1}^{\gamma/\alpha}}=\frac{R^{\frac{d}{d_{\mathrm{c},\alpha}}}}{c_{1}^{\gamma/\alpha}}\gg R.$ (18)
Therefore we can apply (10) and Proposition 3 at $\beta_{*}$. Together, they give
[…]
uniformly in $R$ large and in $x\in\mathbb{T}_{R}^{d}$.
The plateau term is
$\frac{\chi(\beta_{*})}{V}\asymp\frac{1}{Vt_{*}^{\gamma}}=\frac{V^{2/d_{\mathrm{c}}}}{Vc_{1}^{\gamma}}=\frac{1}{c_{1}^{\gamma}V^{1-\frac{2}{d_{\mathrm{c}}}}},$ (20)
as desired. Finally, we sum $G_{R,\beta_{*}}(x)$ over $\mathbb{T}_{R}^{d}$ to get
[…]
since $V^{\frac{2}{d_{\mathrm{c}}}}=R^{\frac{2d}{d_{\mathrm{c}}}}$ dominates $R^{\alpha}$ when $d>d_{\mathrm{c},\alpha}=\frac{\alpha}{2}d_{\mathrm{c}}$. This completes the proof. ∎
IV Long-range self-avoiding walk
In this section we verify Hypotheses 1 and 2 for the spread-out LR self-avoiding walk (SAW) with $\alpha\in(0,2)$ in dimensions $d>d_{\mathrm{c},\alpha}=2\alpha$. This proves that Theorem 1 applies in this setting, with $\gamma=1$.
The self-avoiding walk model is defined as follows. For infinite-volume, we fix a lattice-symmetric, summable kernel $J:{\mathbb{Z}^{d}}\to[0,\infty)$, and we write $J_{x,y}=J(y-x)$ for all $x,y\in{\mathbb{Z}^{d}}$. A self-avoiding walk is a finite path $\omega=(\omega(0),\omega(1),\ldots,\omega(|\omega|))$ consisting of $|\omega|$ steps, with each $\omega(i)$ in $\mathbb{Z}^{d}$ and with $\omega(i)\neq\omega(j)$ when $i\neq j$. The two-point function is defined by
$G_{\beta}(x)=\sum_{\omega:0\to x}\beta^{|\omega|}\prod_{i=1}^{|\omega|}J_{\omega(i-1),\omega(i)},$ (22)
where the sum is over all self-avoiding walks from 0 to $x$ of any length (including the zero-step walk when $x=0$). The torus two-point function $G_{R,\beta}(x)$ is defined similarly, with the sum over self-avoiding walks on the torus and with $J_{x,y}$ replaced by the periodised kernel
$\bar{J}_{R;x,y}=\sum_{u\in\mathbb{Z}^{d}}J_{x,y+Ru}\qquad(x,y\in\mathbb{T}_{R}^{d}).$ (23)
The spread-out LR model is defined by choosing $J$ to have the form
$J_{0,x}\asymp\frac{1}{L^{d}}\Big(\frac{1}{1+|x|/L}\Big)^{d+\alpha},$ (24)
where $L$ is a large parameter. The reciprocal $L^{-1}$ provides a small parameter which permits expansion methods to be used, specifically the lace expansion [33]. In the following, we assume that $L$ is large enough, and verify Hypotheses 1 and 2 for the spread-out long-range SAW.
Verification of Hypothesis 1..
[…]
was proved recently [19], once more using the lace expansion. These substantial results establish (6) with $g=g_{\mathrm{LR}}$ (with $c=1$) and $\xi(\beta)=L({\beta_{\mathrm{c}}}-\beta)^{-1/\alpha}$, which does satisfy $\xi\asymp t^{-\nu}$ with $\nu=1/\alpha=\gamma/\alpha$.
It remains to verify (7). We use the convolution $(f_{1}*f_{2})(x)=\sum_{y\in\mathbb{Z}^{d}}f_{1}(x-y)f_{2}(y)$. Since $G_{\beta}(0)=1$, we only need to consider $x\neq 0$. For this, we begin with the differential inequality
$\beta\frac{\mathrm{d}}{\mathrm{d}\beta}G_{\beta}(x)\leq(G_{\beta}*G_{\beta})(x).$ (26)
This is proved by observing that the factor $|\omega|$ brought down by differentiating $\beta^{|\omega|}$ in (22) may be regarded as a sum over nonzero points in the walk $\omega$. Given a point on the walk, neglecting the self-avoidance between the parts of the walk before and after that point gives an upper bound, which results in (26) (see [34, Lemma 1.5.2] for details).
[…]
Integration of the above from $\beta$ to $\beta_{\mathrm{c}}$}, together with the power-law decay of the critical two-point function $G_{{\beta_{\mathrm{c}}}}$, gives
$G_{\beta}(x)$ $=G_{\beta_{\mathrm{c}}}(x)-(G_{\beta_{\mathrm{c}}}(x)-G_{\beta}(x))$
[…]
for some constants $a,A>0$. Since $\xi(\beta)\asymp({\beta_{\mathrm{c}}}-\beta)^{-1/\alpha}$, the subtracted term is at most
[…]
with some constant $A^{\prime}$, and this is negligible compared to the main term $a|x|^{-(d-\alpha)}$ if $|x|\!|\!|\leq\varepsilon\xi(\beta)$ with an $\varepsilon>0$ sufficiently small. This verifies (7) with $s_{1}=\varepsilon$. ∎
Verification of Hypothesis 2..
Let $\pi:{\mathbb{Z}^{d}}\to\mathbb{T}_{R}^{d}$ denote the natural projection, which we also apply to $\omega$ component-wise. We first prove the upper bound, which asserts that
$G_{R,\beta}(x)$ $=\sum_{\omega_{R}:0\to x}\beta^{|\omega_{R}|}\prod_{j=1}^{|\omega_{R}|}\bar{J}_{R;\omega_{R}(j-1),\omega_{R}(j)}$
$\leq\sum_{\begin{subarray}{c}x^{\prime}\in\mathbb{Z}^{d}\\ \pi(x^{\prime})=x\end{subarray}}\sum_{\omega:0\to x^{\prime}}\beta^{|\omega|}\prod_{j=1}^{|\omega|}J_{\omega(j-1),\omega(j)},$ (30)
where the first sum is over torus SAWs $\omega_{R}$ and the second line involves a sum over $\mathbb{Z}^{d}$} SAWs $\omega$.
By the definition of the torus kernel $\bar{J}_{R}$ in (23), the torus product can be unwrapped to give
$\prod_{j=1}^{|{\omega_{R}}|}\bar{J}_{R;{\omega_{R}}(j-1),{\omega_{R}}(j)}=\sum_{\omega}\mathds{1}_{\{\pi(\omega)={\omega_{R}}\}}\prod_{j=1}^{|\omega|}J_{\omega(j-1)\omega(j)},$ (31)
where the indicator function $\mathds{1}_{\{\pi(\omega)={\omega_{R}}\}}$ is $1$ if $\omega$ projects to $\omega_{R}$ and otherwise is 0. We multiply by $\beta^{|{\omega_{R}}|}=\beta^{|\omega|}$ and sum over $\omega_{R}$} from 0 to $x$. The walk $\omega$ must end at some $x^{\prime}\in{\mathbb{Z}^{d}}$ with $\pi(x^{\prime})=x$, so we get
$G_{R,\beta}(x)=\sum_{\begin{subarray}{c}x^{\prime},\,\omega:0\to x^{\prime}\\ \pi(x^{\prime})=x\end{subarray}}\mathds{1}_{\{\pi(\omega)\text{ is SA}\}}\beta^{|\omega|}\prod_{j=1}^{|\omega|}J_{\omega(j-1)\omega(j)},$ (32)
where SA denotes self-avoiding. Since $\pi(\omega)$ is SA implies that $\omega$ is SA, by relaxing the indicator function we get
$G_{R,\beta}(x)$ $\leq\sum_{\begin{subarray}{c}x^{\prime},\omega:0\to x^{\prime}\\ \pi(x^{\prime})=x\end{subarray}}\mathds{1}_{\{\omega\text{ is SA}\}}\beta^{|\omega|}\prod_{j=1}^{|\omega|}J_{\omega(j-1)\omega(j)}$
$=\sum_{x^{\prime}:\pi(x^{\prime})=x}G_{\beta}(x^{\prime})=\Gamma_{R,\beta}(x).$ (33)
This proves the upper bound of (10).
For the lower bound of (10), our goal is to find constants $c_{0},s_{2}>0$ such that if $\xi(\beta)\geq s_{2}R$ then
$\Gamma_{R,\beta}(x)-G_{R,\beta}(x)\leq c_{0}\frac{\chi(\beta)^{d_{\mathrm{c}}/2}}{V}\Gamma_{R,\beta}(x).$ (34)
To begin, we observe from (IV) that the difference $\Gamma_{R,\beta}-G_{R,\beta}$ arises from self-avoiding walks $\omega$ whose projection $\pi(\omega)$ is not self-avoiding. Such walks must visit two distinct points $y,y^{\prime}$ in $\mathbb{Z}^{d}$} with $\pi(y)=\pi(y^{\prime})$. Diagrammatically, this is  [Uncaptioned image]  where the hollow circles are $0,x^{\prime}$ and the filled squares are $y,y^{\prime}$. By neglecting the mutual avoidance of the three diagram lines, the difference $\Gamma_{R,\beta}(x)-G_{R,\beta}(x)$ is therefore bounded above by
$\sum_{y\in\mathbb{Z}^{d}}G_{\beta}(y)\sum_{y^{\prime}\neq y:\pi(y^{\prime})=y,}G_{\beta}(y^{\prime}-y)\sum_{x^{\prime}:\pi(x^{\prime})=x}G_{\beta}(x^{\prime}-y^{\prime}).$ (35)
The last sum over $x^{\prime}$ is exactly $\Gamma_{R,\beta}(x-\pi(y))$. Then we can use (25) and Lemma 2 to bound the middle sum over $y^{\prime}$ by $O(\xi^{\alpha}/V)$. The remaining sum over $y$ is
$\sum_{y\in{\mathbb{Z}^{d}}}G_{\beta}(y)\Gamma_{R,\beta}(x-\pi(y))=\sum_{u\in{\mathbb{Z}^{d}}}(G_{\beta}*G_{\beta})(x+Ru).$ (36)
[…]
To conclude, we take $s_{2}=\max\{\frac{1}{2},\frac{3}{2s_{1}}\}$, and use $|x|\!|\!|^{\alpha}\leq(\frac{R}{2})^{\alpha}\leq(s_{2}R)^{\alpha}\leq\xi^{\alpha}$. Since also $\xi^{\alpha}\asymp\chi$, we find that there is a constant $c_{0}^{\prime}$ such that
[…]
By Proposition 3, the above inequality gives the desired result (34), since $d_{\mathrm{c}}=4$. ∎
V Further applications of the plateau theorem
The plateau theorem has been isolated here, and also applied to long-range models, for the first time. However, it has previously been used implicitly in multiple contexts. We now summarise the relevant literature; the conclusions are presented in Table 1. (Logarithmic corrections at the critical dimension are discussed in [30, 31, 35, 36, 37] for SR models.)
SAW/Spin Percolation BP
$d_{\mathrm{c}}$ 44 66 88
$d_{\mathrm{c},\alpha}$ $2\alpha$ $3\alpha$ $4\alpha$
$\gamma$ 11 11 $1/2$
window $V^{-1/2}$ $V^{-1/3}$ $V^{-1/2}$
$\chi_{R}$ $V^{1/2}$ $V^{1/3}$ $V^{1/4}$
plateau $V^{-1/2}$ $V^{-2/3}$ $V^{-3/4}$
Table 1: Finite-size scaling exponents for short-range models with $d>d_{\mathrm{c}}$ and long-range models with $d>d_{\mathrm{c},\alpha}=\frac{\alpha}{2}d_{\mathrm{c}}$. Rigorous results for long-range models are partial, as discussed in the text.
The results concern the following symmetric kernels $J$ on $\mathbb{Z}^{d}$ and $\bar{J}_{R}$ on $\mathbb{T}_{R}^{d}$ for short-range and long-range models:
  • SR: $J_{x,y}=1$ for nearest-neighbour $x,y$, and otherwise $J_{x,y}=0$.
  • spread-out SR: $J_{x,y}=L^{-d}$ for $0<|x-y|\leq L$, and otherwise $J_{x,y}=0$ ($L$ is fixed and large);
  • LR: $J_{x,y}\asymp|x-y|^{-(d+\alpha)}$ with $\alpha\in(0,2)$;
  • spread-out LR: $J_{x,y}\asymp L^{-d}(1+|x-y|/L)^{-(d+\alpha)}$ with $\alpha\in(0,2)$ for $d>1$ and $\alpha\in(0,1)$ for $d=1$ ($L$ is fixed and large).
  • torus versions: $\bar{J}_{R;x,y}=\sum_{u\in\mathbb{Z}^{d}}J_{x,y+Ru}$.
Self-avoiding walk ($d_{\mathrm{c}}=4$). The self-avoiding walk on $\mathbb{Z}^{d}$ has two-point function defined by (22), with any of the above choices of $J$. The torus two-point function is defined with a sum over torus walks and with $J$ replaced by $\bar{J}_{R}$. Theorem 1 has been proved in the following settings, using the lace expansion:
  • SR (nearest-neighbour) SAW for $d>4$ [14, 15];
  • spread-out LR SAW for $d>2\alpha$ in Section IV.
Branched Polymer (BP) models ($d_{\mathrm{c}}=8$). Let $\mathcal{C}$ denote either a lattice animal (finite connected bond cluster) or a lattice tree (acyclic animal), and let $|\mathcal{C}|$ denote the number of bonds in $\mathcal{C}$. The infinite-volume two-point function is defined by
$G_{\beta}(x)=\sum_{\mathcal{C}\ni 0,x}\beta^{|\mathcal{C}|}\prod_{\{x,y\}\in\mathcal{C}}J_{x,y},$ (39)
with the product over all bonds in $\mathcal{C}$. The torus two-point function $G_{R,\beta}(x)$ is defined by the sum over $\mathcal{C}$ in the torus, with $J$ replaced by $\bar{J}_{R}$. Theorem 1 has been proved for:
  • Spread-out SR BP for $d>8$ [18].
Percolation ($d_{\mathrm{c}}=6$). Percolation is a probabilistic model in which each bond $b$ is occupied with probability $1-e^{-\beta J_{b}}$ and vacant with probability $e^{-\beta J_{b}}$. Bond occupations are independent. A configuration of occupied bonds forms a subgraph of the complete graph on $\mathbb{Z}^{d}$ (or $\mathbb{T}_{R}^{d}$), and we say $0\leftrightarrow x$ if 0 and $x$ are in the same connected component of this subgraph. The two-point functions $G_{\beta}(x),G_{R,\beta}(x)$ are defined by $\mathrm{Prob}_{\beta}(0\leftrightarrow x)$ on $\mathbb{Z}^{d}$ or on the torus, respectively.
  • Theorem 1 has been proved for SR percolation for $d\geq 11$, and for spread-out SR percolation for $d>6$. Moreover, both asymptotic relations in (11) have been proved throughout the entire critical window $|t|\lesssim V^{-1/3}$, including the critical point and the supercritical regime [16]. The verification of Hypothesis 2 is achieved via an exploration process which couples percolation on $\mathbb{Z}^{d}$ and the torus [38].
  • For spread-out LR percolation, the near-critical upper bound (6) is proved in [19] for $\alpha\in(0,2)$. It is also proved in [39] for $\alpha\in(0,1)$ for LR percolation without requiring spread-out. To extend this to prove Theorem 1 for LR percolation would require an adaptation of the above-mentioned coupling to the LR setting.
Spin systems ($d_{\mathrm{c}}=4$). The Ising model has spins $\varphi_{x}=\pm 1$ with the Hamiltonian $H=-\sum_{x\sim y}J_{x,y}\varphi_{x}\varphi_{y}$ for $\mathbb{Z}^{d}$, and with instead $\bar{J}_{R}$ for the torus. We also consider the $|\varphi|^{4}$ model with $n$-component continuous spins $\varphi_{x}\in\mathbb{R}^{n}$; its Hamiltonian is discussed in Appendix D. The two-point functions $G_{\beta}(x),G_{R,\beta}(x)$ are defined by the expectation $\frac{1}{n}\langle\varphi_{0}\cdot\varphi_{x}\rangle_{\beta}$. For the Ising model, $\beta$ is the inverse temperature. For the $|\varphi|^{4}$ model, the role of $\beta$ is played by $\nu$ where $\nu$ is the quadratic coupling constant (see Appendix D).
  • Theorem 1 has been proved for the SR Ising model for $d>4$ [17]. The proof relies on the fact that (6)–(7) are proved in [40]. The verification of Hypothesis 2 uses a coupled exploration process and the random current representation of the Ising model [41].
  • For the spread-out LR Ising model, the upper bound (6) is proved in [19]. To extend this to prove Theorem 1 would require adaptation of the above-mentioned coupled exploration process to the LR setting.
  • For the SR $1$-component $\varphi^{4}$ model, (6)–(7) are proved in [40] for $d>4$. For the torus, a rigorous renormalisation group method has recently been used to prove that for $d>4$ the bounds of Theorem 1 hold exactly at the infinite-volume critical point for both $\chi_{R}(\beta_{c})$ and $G_{R,\beta_{c}}$ [37], and with logarithmic corrections for $d=4$. Stronger results have been proved on the hierarchical lattice for $d\geq 4$ [30, 31]. We discuss this in more detail in Section VI.
Some ideas behind the proofs of the lower bound of (10) for BP, percolation, and the Ising model are sketched in Appendix B.
VI Universal profiles
Theorem 1 shows that the torus two-point function has a plateau provided Hypotheses 1 and 2 are satisfied. However, it does not provide precise information on the behaviour of the two-point function or susceptibility within the critical window. We now present two conjectures on the behaviour of the torus susceptibility and plateau throughout the critical window. We parametrise the window using
${\beta_{\mathrm{c}}}(s)={\beta_{\mathrm{c}}}+sa_{d}V^{-\frac{2}{\gamma d_{\mathrm{c}}}},$ (40)
for $s\in\mathbb{R}$, for some non-universal positive constant $a_{d}$, and with $\beta_{\mathrm{c}}$} the infinite-volume critical point. We write $f\sim g$ to mean $\lim f/g=1$.
Conjecture 4.
Let $d>d_{\mathrm{c},\alpha}$ and $s\in\mathbb{R}$. For each of the models (SAW, Ising, $|\varphi|^{4}$, percolation, BP), there is a model-dependent profile function $f:\mathbb{R}\to(0,\infty)$ (the same profile for SR and LR) and positive constants $a_{d},b_{d}$, such that, as $R\to\infty$,
$\chi_{R}({\beta_{\mathrm{c}}}(s))\sim b_{d}f(s)V^{\frac{2}{d_{\mathrm{c}}}}.$ (41)
Also, as $R\to\infty$, for every $x\in\mathbb{T}_{R}^{d}$,
$G_{R,{\beta_{\mathrm{c}}}(s)}(x)\sim G_{{\beta_{\mathrm{c}}}}(x)+\frac{b_{d}f(s)}{V^{1-\frac{2}{d_{\mathrm{c}}}}}.$ (42)
For k>1k>-1 and $s\in\mathbb{R}$, let
$I_{k}(s)=\int_{0}^{\infty}x^{k}e^{-\frac{1}{4}x^{4}-\frac{1}{2}sx^{2}}\mathrm{d}x.$ (43)
Also, we define
$f_{\rm perc}(s)=\int_{0}^{\infty}\frac{\Psi(x^{3/2})}{\sqrt{2\pi x}}e^{-\frac{1}{6}x^{3}+\frac{1}{2}sx^{2}-\frac{1}{2}s^{2}x}\mathrm{d}x,$ (44)
where $\Psi(z)=\mathbb{E}\exp\{z\int_{0}^{1}W^{*}(t)\mathrm{d}t\}$ is the moment generating function of the Brownian excursion area.
Conjecture 5.
The profile functions $f(s)$ are given by the functions in Table 2, for both SR and LR models for all $d\geq d_{\mathrm{c},\alpha}$ (including the upper critical dimension).
SAW Spin ($n\geq 1$ components) Percolation BP
$I_{1}$ $I_{n+1}/(nI_{n-1})$ $f_{\mathrm{perc}}$ $I_{0}$
Table 2: Conjectured universal profile for short-range models with $d\geq d_{\mathrm{c}}$ and long-range models with $d\geq d_{\mathrm{c},\alpha}$.
For the $n$-component $|\varphi|^{4}$ model with $d>4$, the profile
$f_{n}(s)=\frac{I_{n+1}(s)}{nI_{n-1}(s)}$ (45)
is computed in [28, 29]. That computation uses a renormalisation group (RG) analysis which we recall in Appendix D. A rigorous RG analysis confirms this profile $f_{n}$ for the weakly-coupled $n$-component $|\varphi|^{4}$ model on a hierarchical lattice [30, 31], including for $d=d_{\mathrm{c}}=4$ with logarithmic corrections (see Appendix C). More generally, at the upper critical dimension $d_{\mathrm{c},\alpha}$, we expect Conjecture 4 to continue to hold with the same profile function $f(s)$, but with logarithmic corrections in the window, susceptibility, and plateau scales.
Despite the fact that $f_{n}$ was first computed as long ago as 1985, the existence and form of profile functions in the scaling window have received scant attention in the finite-size scaling literature—for example it is not explicitly mentioned in any of [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13]. One of our goals in this paper is to draw renewed attention to $f_{n}$ and other profiles.
Some of the conjectured profiles have been proven to apply to models defined on the complete graph ($V$ nodes with edges connecting each pair of nodes). The $n=1$ profile $f_{1}$ agrees with the profile for the Curie–Weiss model (Ising model on complete graph) [42]. The common profile function for lattice trees and lattice animals has been proven to be $I_{0}$ on the complete graph [43]. Integration by parts in the denominator of (45) leads to the formula
$f_{n}(s)=\frac{I_{n+1}(s)}{I_{n+3}(s)+sI_{n+1}(s)},$ (46)
which extends the definition of $f_{n}$ to all n>2n>-2. The $n=0$ case reduces to $f_{0}(s)=I_{1}(s)$ (since $I_{3}(x)+sI_{1}(s)=1$) as in Table 2; this is the profile for SAW on the complete graph [30, Appendix B] and is consistent with SAW being the $n=0$ spin model [44]. The limiting case $f_{-2}(s)=s^{-1}$ is the correct profile for a Gaussian model. Percolation on the complete graph is the Erdős–Rényi random graph. Its profile function has been proved to be $f_{\mathrm{perc}}$ in [45, Corollary 3.9] (see [17, Appendix C] for the identification of the profile function with $f_{\mathrm{perc}}$). At the moment of final revision of our paper, equation (41) in Conjecture 4 has been proved for percolation on the high-dimensional torus [46, Theorem 2.1] for SR spread-out models in dimensions $d>6$, and for the nearest-neighbour model for $d$ much greater than $6$.
VII Free boundary conditions
Models with free boundary conditions are formulated in a discrete $d$-dimensional box $\Lambda_{R,d}$ of side $R$ and volume $V=R^{d}$. They are defined by restricting a kernel $J_{x,y}$ on $\mathbb{Z}^{d}$ to satisfy $J_{x,y}=0$ if $x$ or $y$ is not in the box $\Lambda_{R,d}$. We use the superscript $\mathrm{F}$ to indicate FBC. Unlike periodic boundary conditions (the torus) there is no wrapping.
For FBC, the finite-size scaling at the infinite-volume critical point is conjectured—and in some cases proved—to be different from PBC, namely that in dimensions $d>d_{\mathrm{c},\alpha}$, the susceptibility and two-point function obey
$\chi_{R}^{\mathrm{F}}({\beta_{\mathrm{c}}})\asymp R^{\alpha},\qquad G_{R,{\beta_{\mathrm{c}}}}^{\mathrm{F}}(0,x)\asymp\frac{1}{|x|^{d-\alpha}},$ (47)
with $x$ away from the box boundary for the latter. (Since translation invariance is broken, we write both points in $G_{R,{\beta_{\mathrm{c}}}}^{\mathrm{F}}(0,x)$; the origin is the centre of the box.) In contrast with (4) for PBC, there is no plateau in (47). This has been proved for the SR Ising model [47] (with the hypothesis supplied by [40, 22]) and SR percolation [48].
We conjecture that the $V^{-\frac{2}{\gamma d_{\mathrm{c}}}},V^{\frac{2}{d_{\mathrm{c}}}},V^{-(1-\frac{2}{d_{\mathrm{c}}})}$ scales of window, susceptibility, and plateau, and also the profile functions of Table 2, are exactly duplicated at a pseudocritical point $\beta_{R,\mathrm{c}}$ shifted into the ordered phase as $\beta_{R,\mathrm{c}}={\beta_{\mathrm{c}}}+\text{const}R^{-1/\nu}$. (This shift has been observed previously by other authors, e.g., [4].) Explicitly, with a window around $\beta_{R,\mathrm{c}}$ parametrised by
$\beta_{R,\mathrm{c}}(s)=\beta_{R,\mathrm{c}}+sa_{d}V^{-\frac{2}{\gamma d_{\mathrm{c}}}},$ (48)
we have the following conjecture. Its constants $a_{d}$ and $b_{d}$ are not the same as for PBC. The restriction $|x|/R\to 0$ for (50) is present to account for possible $x$ dependence in the profile for $x$ on the scale of $R$. As noted in (5), the profile dominates the Gaussian decay as soon as $|x|\gg R^{p}$ with $p=\frac{d-\alpha\frac{d}{d_{\mathrm{c},\alpha}}}{d-\alpha}<1$, so the restriction $|x|/R\to 0$ still includes a large part of the box where the universal profile would apply.
Conjecture 6.
Let $d>d_{\mathrm{c},\alpha}$ and $s\in\mathbb{R}$. For each of the models (SAW, Ising, $|\varphi|^{4}$, percolation, BP), there is a pseudocritical point $\beta_{R,\mathrm{c}}={\beta_{\mathrm{c}}}+{\rm const}R^{-1/\nu}$ such that, as $R\to\infty$,
$\chi_{R}^{\mathrm{F}}(\beta_{R,\mathrm{c}}(s))\sim b_{d}f(s)V^{\frac{2}{d_{\mathrm{c}}}}$ (49)
with the same profile function $f$ from Table 2. Also, as $R\to\infty$, for $x\in\Lambda_{R,d}$ with $|x|/R\to 0$,
$G^{\mathrm{F}}_{R,\beta_{R,\mathrm{c}}(s)}(0,x)\sim G_{{\beta_{\mathrm{c}}}}(x)+\frac{b_{d}f(s)}{V^{1-\frac{2}{d_{\mathrm{c}}}}}.$ (50)
Since $d>d_{\mathrm{c},\alpha}=\frac{\alpha}{2}d_{\mathrm{c}}$, the shift $R^{-\frac{1}{\nu}}=V^{-\frac{\alpha}{\gamma d}}$ is larger than the window width $V^{-\frac{2}{\gamma d_{\mathrm{c}}}}$, and the PBC and FBC windows do not overlap. This shifting of the PBC scaling behaviour to the FBC pseudocritical point was observed numerically for the 5D SR Ising model and is discussed in [9, 4] (contrary to the conclusions of [13]).
Conjecture 6, and the Gaussian scaling $\chi_{R}^{\mathrm{F}}({\beta_{\mathrm{c}}})\asymp R^{2}$ at the infinite-volume critical point, have been proved for the hierarchical $n$-component $|\varphi|^{4}$ model in [30, 31]. This is true also for $d=d_{\mathrm{c}}=4$ with a logarithmic correction to the shift, which instead of $R^{-1/\nu}=R^{-1/2}$ becomes $R^{-1/2}(\log R)^{\frac{n+2}{n+8}}$. In general, for all models discussed above, at their upper critical dimension we expect logarithmic corrections to the window, the susceptibility, and the plateau to be identical for the (non-overlapping) FBC and PBC windows.
VIII Conclusion
We have developed a unified theory for finite-size scaling under periodic boundary conditions above the upper critical dimension, and have demonstrated that polymer models, percolation, and spin systems all fit into this general framework. The theory computes exponents for the scaling window, the critical susceptibility, and the two-point function plateau. In our work, a larger-than-system correlation length occurs as the correlation length of an unwrapped model. We propose conjectures for the precise behaviour of the models inside the critical window, in which the susceptibility and the torus plateau are governed by universal profiles. These profiles are computed on the complete graph and for a hierarchical model, and are conjectured to also apply to short- and long-range models with PBC. Under free boundary conditions, the same behaviour and the same universal profiles are conjectured to apply in a shifted critical window around a pseudocritical point.
Acknowledgements
We thank our collaborators on aspects of this work: Tom Hutchcroft, Emmanuel Michta, Romain Panis. The work of YL and GS was supported in part by NSERC of Canada. YL and GS thank the Isaac Newton Institute for Mathematical Sciences, Cambridge, for support and hospitality during the programme Stochastic systems for anomalous diffusion, where work on this paper was undertaken; this work was supported by EPSRC grant EP/R014604/1. The work of GS was supported in part by the Clay Mathematics Institute. JP thanks the Seoul National University for hospitality during a visit when this work was partially undertaken. JP was supported by Basic Science Research Program through the National Research Foundation of Korea funded by the Ministry of Science and ICT (RS2025-00518980).
Appendix A Elementary lemmas
This appendix contains two lemmas. The first is a restatement of Lemma 2, with proof. The proof uses the simple observation that for $x\in\mathbb{T}_{R}^{d}$ (identified with a point in $\mathbb{Z}^{d}$ with $|x|\leq R/2$) and for nonzero $u\in{\mathbb{Z}^{d}}$, the points $x+Ru$ and $u$ are simply related by
$|x+Ru|$ $\geq R|u|-\frac{R}{2}\geq\frac{R}{2}|u|,$ (51)
$|x+Ru|$ $\leq R|u|+\frac{R}{2}\leq\frac{3}{2}R|u|.$
Lemma 7.
Let $d\geq 1$, $R\geq 1$, $a>0$ and $\xi>0$. Suppose $g:[0,\infty)\to[0,\infty)$ satisfies $g(s)\leq(1+s)^{-(a+\varepsilon)}$ for some $\varepsilon>0$. Then there is a constant $C=C(d,a,\varepsilon)>0$ such that
$\sum_{u\in\mathbb{Z}^{d}:u\neq 0}\frac{1}{|x+Ru|^{d-a}}g\bigg(\frac{|x+Ru|}{\xi}\bigg)\leq C\frac{\xi^{a}}{R^{d}}$
for all $x\in{\mathbb{Z}^{d}}$ with $|x|\leq R/2$.
Proof.
We may assume that $g(s)=(1+s)^{-(a+\varepsilon)}$. By (51) and the monotonicity of $g$ it is enough to obtain an upper bound of order $\xi^{a}/R^{d}$ for
$\sum_{u\in\mathbb{Z}^{d}:u\neq 0}\frac{1}{(\frac{R}{2}|u|)^{d-a}}g\bigg(\frac{R}{2\xi}|u|\bigg)$
$\qquad\lesssim\frac{1}{R^{d-a}}\sum_{N=1}^{\infty}N^{a-1}g\bigg(\frac{R}{2\xi}N\bigg).$ (52)
If $0<a\leq 1$, the summand is decreasing in $N$, so we can bound the sum by the integral
$\int_{0}^{\infty}t^{a-1}g\bigg(\frac{R}{2\xi}t\bigg)\mathrm{d}t$ $=\bigg(\frac{2\xi}{R}\bigg)^{a}\int_{0}^{\infty}y^{a-1}g(y)\mathrm{d}y$
$=C_{a,\varepsilon}\frac{\xi^{a}}{R^{a}},$ (53)
using $a>0$ and $\varepsilon>0$.
If $a>1$, by the definition of $g$ it suffices to bound
$\sum_{N=1}^{\lfloor 2\xi/R\rfloor}N^{a-1}+\sum_{N=\max\{\lceil 2\xi/R\rceil,1\}}^{\infty}N^{a-1}\bigg(\frac{R}{2\xi}N\bigg)^{-(a+\varepsilon)}.$ (54)
The first term is $O(\xi^{a}/R^{a})$ since $a>1$. The second term is a multiple of
$\frac{\xi^{a+\varepsilon}}{R^{a+\varepsilon}}\sum_{N=\max\{\lceil 2\xi/R\rceil,1\}}^{\infty}N^{-1-\varepsilon}.$ (55)
If the maximum in the lower summation limit is $1$ then $\xi/R\leq 1$ and ($5$) is of order $(\xi/R)^{a+\varepsilon}\leq(\xi/R)^{a}$. If instead the maximum is $\lceil 2\xi/R\rceil$ then (55) is bounded by a multiple of $(\xi/R)^{a+\varepsilon}(\xi/R)^{-\varepsilon}=(\xi/R)^{a}$. This completes the proof. ∎
The second lemma is an elementary convolution estimate, which is applied in Section IV to verify Hypothesis 2 for the spread-out LR SAW.
Lemma 8.
Let $\alpha>0$, $d>2\alpha$, and $\xi\in[1,\infty]$. Suppose $G:{\mathbb{Z}^{d}}\to[0,\infty)$ obeys
[…]
Then there is a constant independent of $\xi$ such that
[…]
Proof.
[…]
$(G*G)(x)\lesssim\frac{\xi^{3\alpha}}{|x|^{d+\alpha}}.$ (58)
In the sum $(G*G)(x)=\sum_{y\in{\mathbb{Z}^{d}}}G(y)G(x-y)$, either |xy||x-y| or $|y|$ must be at least $\frac{1}{2}|x|$. In the first case, we use the assumed bound on G(xy)G(x-y) to get
$\sum_{y\in{\mathbb{Z}^{d}}:|x-y|\geq\frac{1}{2}|x|}G(y)G(x-y)\lesssim\frac{\xi^{2\alpha}}{(\frac{1}{2}|x|)^{d+\alpha}}\sum_{y\in{\mathbb{Z}^{d}}}G(y),$ (59)
and we use the hypothesis again to see that
[…]
This gives an upper bound $\xi^{3\alpha}/|x|^{d+\alpha}$ for the $|x-y|\geq\frac{1}{2}|x|$ part of the convolution. The $|y|\geq\frac{1}{2}|x|$ part is analogous, with the decay coming from $G(y)$ instead of G(xy)G(x-y). This gives (58) and completes the proof. ∎
Appendix B Lower bound of Hypothesis 2 for short-range models
Hypothesis 1 is proved for various models in [14, 50, 51, 16, 40, 18, 15, 39, 19]. In this appendix, we briefly sketch the proof of the lower bound of (10) in Hypothesis 2, for short-range self-avoiding walk, branched polymers, percolation, and the Ising model. References are given where full details can be found.
We use the fact that Hypothesis 1 has been established in previous works with $g=g_{\mathrm{SR}}$. Our goal is to prove that there are constants $c_{0},s_{2}$ such that, when $\xi(\beta)\geq s_{2}R$, we have
$\Gamma_{R,\beta}(x)-G_{R,\beta}(x)\leq c_{0}\frac{\chi(\beta)^{d_{\mathrm{c}}/2}}{V}\Gamma_{R,\beta}(x).$ (61)
The model dependence in this bound arises via the diagrams for $\Gamma_{R,\beta}(x)-G_{R,\beta}(x)$ depicted in Figure 2, which project onto the torus diagrams in Figure 3. We estimate the diagrams using the torus convolution
$(f\star g)(x)=\sum_{y\in\mathbb{T}_{R}^{d}}f(x-y)g(y)$ (62)
for functions $f,g:\mathbb{T}_{R}^{d}\to\mathbb{R}$. From the near-critical upper bound (6), a convolution estimate, and Lemma 2, the $m$-fold convolution of $\Gamma_{R,\beta}$ with itself satisfies, for $d>2m$,
[…]
Together with $|x|^{2}\leq(\frac{R}{2})^{2}\lesssim\xi^{2}\asymp\chi$ and the lower bound of $\Gamma_{R,\beta}(x)$ from Proposition 3, this implies the useful inequality
$\Gamma_{R,\beta}^{\star m}(x)\lesssim\chi^{m-1}\Gamma_{R,\beta}(x).$ (64)
Figure 2: $\mathbb{Z}^{d}$ configurations contributing to $\Gamma_{R,\beta}(x)-G_{R,\beta}(x)$. First line, left to right: SAW, Ising, percolation. Second line: two topologies for BP. Thin lines represent $G_{\beta}$; the bold line for Ising represents $\Gamma_{R,\beta}$. Hollow vertices are 0 and $x^{\prime}=x+Ru$, filled vertices are summed over $\mathbb{Z}^{d}$, and box vertices are summed over torus equivalent, distinct, unordered pairs in $\mathbb{Z}^{d}$.
Figure 3: Torus configurations contributing to $\Gamma_{R,\beta}(x)-G_{R,\beta}(x)$. First line, left to right: SAW, Ising, percolation. Second line: three diagrams for BP. Solid lines represent $\Gamma_{R,\beta}$, dashed lines represent $\chi/V$, two dashed lines connected by $\times$ represent $\chi^{2}/V$. Hollow vertices are $0,x$ and filled vertices are summed over the torus.
B.1 Self-avoiding walk
For SR SAW, the inequality (61) follows exactly as in the derivation of (10) for LR SAW in Section IV, with $\alpha$ set equal to $2$. The SAW diagram of Figure 2 was encountered in Section IV. It projects to the torus SAW diagram in Figure 3.
B.2 Branched polymers
Similarly to SAW, the difference $\Gamma_{R,\beta}(x)-G_{R,\beta}(x)$ for lattice trees arises from $\mathbb{Z}^{d}$ trees that do not project to a torus tree. These trees must contain two distinct points that project to the same torus point. This can happen with two topologies, as depicted in Figure 2, depending on whether the unique path joining the distinct points intersects the path joining 0 and $x^{\prime}$. Analogous considerations to those for SAW lead to an upper bound by the three BP diagrams in Figure 3 [18]. Using (64), these diagrams are bounded, for $m=4,3,2$ respectively, by
$\frac{\chi^{5-m}}{V}\Gamma_{R,\beta}^{\star m}(x)$ $\lesssim\frac{\chi^{5-m}}{V}\chi^{m-1}\Gamma_{R,\beta}(x)=\frac{\chi^{4}}{V}\Gamma_{R,\beta}(x).$ (65)
The desired inequality (61) then follows since $d_{\mathrm{c}}=8$.
For lattice animals, minimal additional care is required to specify the unwrapping operation. Once this is done, the rest is the same as for lattice trees [18].
B.3 Percolation
A coupling of percolation on the torus and the infinite lattice $\mathbb{Z}^{d}$ is discussed in detail in [38, 16]. It involves an exploration process which provides the unwrapping of a torus percolation cluster to the infinite lattice. From this, the upper bound $G_{R,\beta}(x)\leq\Gamma_{R,\beta}(x)$ follows directly. It also leads to an upper bound on the difference $\Gamma_{R,\beta}(x)-G_{R,\beta}(x)$ in which there is a connection in $\mathbb{Z}^{d}$ from 0 to a point $x^{\prime}$ which projects to $\pi(x^{\prime})=x$, passing through some point $y$, with 0 also connected to a point $y^{\prime}\neq y$ with $\pi(y^{\prime})=\pi(y)$. A generic such configuration is depicted in Figure 2, and its representation as a torus diagram is depicted in Figure 3. By (64) with $m=3$, the torus diagram is bounded by
$\frac{\chi}{V}\Gamma_{R,\beta}^{\star 3}(x)\lesssim\frac{\chi^{3}}{V}\Gamma_{R,\beta}(x).$ (66)
This gives (61) because $d_{\mathrm{c}}=6$.
B.4 Ising model
The unwrapping procedure for the Ising model uses the random current representation, which provides a percolation-like geometric representation for Ising correlation functions. The advantages of this geometric representation were first championed in [41]. A current configuration $\bf n$} is an assignment of a non-negative integer ${\bf n}_{b}$ to each bond $b$. A lattice site $x$ is a source if the sum of ${\bf n}_{b}$ over bonds incident to $x$ is odd. The set of sources of $\bf n$} is denoted $\partial{\bf n}$. Let $w_{\beta}({\bf n})=\prod_{b}\beta^{{\bf n}_{b}}/{\bf n}_{b}!$. The random current representation for the Ising two-point function is
$G_{\beta}(x)=\frac{\sum_{{\bf n}:\partial{\bf n}=\{0,x\}}w_{\beta}({\bf n})}{\sum_{{\bf n}:\partial{\bf n}=\varnothing}w_{\beta}({\bf n})}.$ (67)
Via this representation, a coupling between the Ising model on the torus and on the infinite lattice is established in [17]. The bound $G_{R,\beta}(x)\leq\Gamma_{R,\beta}(x)$ is proved using this coupling. The coupling also shows that $\Gamma_{R,\beta}(x)-G_{R,\beta}(x)$ is dominated by a configuration diagrammatically depicted in Figure 2, which projects to the torus diagram of Figure 3. The latter diagram is bounded using an extension of (64) which implies that
$\frac{\chi}{V}(\Gamma_{R,\beta}\star\Gamma_{R,\beta}^{2}\star\Gamma_{R,\beta})(x)\lesssim\frac{\chi^{2}}{V}\Gamma_{R,\beta}(x).$ (68)
Details are given in [17].
[…]
In our work, a connection with Lee–Yang zeros is not made. On the other hand, the correlation length $\xi$ of the infinite system plays a prominent role. Its value at the point $\beta_{*}$ in the critical window obeys
$\xi(\beta_{*})\asymp(V^{-\frac{2}{\gamma d_{\mathrm{c}}}})^{-\nu}=R^{\frac{2\nu}{\gamma}\frac{d}{d_{\mathrm{c}}}}=R^{\frac{2}{\alpha}\frac{d}{d_{\mathrm{c}}}}=R^{\frac{d}{d_{\mathrm{c},\alpha}}}.$ (69)
The exponent on the right-hand side agrees with and provides an alternate interpretation for ϙ.
For the SR $n$-component $|\varphi|^{4}$ model in the critical dimension $d=d_{\mathrm{c}}=4$, it is well-known and rigorously proved [53, 54] that $\gamma=2\nu=1$ with logarithmic exponents $\hat{\gamma}=2\hat{\nu}=\frac{n+2}{n+8}$. Also, the specific heat has exponent $\alpha=0$ for $n\geq 1$, with $\hat{\alpha}=\frac{4-n}{n+8}$ for $n=1,2,3$, whereas for $n=4$ the specific heat diverges as $\log\log t^{-1}$ and for $n>4$ it does not diverge at all [53]. In particular, $\hat{\alpha}$ is never strictly negative. Let $\hat{\theta}=\frac{4-n}{2(n+8)}=\frac{1}{2}-\hat{\gamma}$, for all $n\geq 1$ including $n\geq 4$. It is argued in [35, (3.6), (4.3)] that under PBC the finite-volume susceptibility has a log correction $\chi_{R}\asymp V^{1/2}(\log V)^{1/2}$ and that the window width is $V^{-1/2}(\log V)^{-\hat{\theta}}$. It was proved rigorously that $\chi_{R}(\beta_{c})\asymp V^{1/2}(\log V)^{1/2}$ in [37] using a rigorous RG analysis [55]. For the hierarchical model, precise asymptotics for both the torus susceptibility and window width were proved in [30].
At the high-temperature edge of the window, the infinite-volume correlation length therefore obeys
$\xi(\beta_{*})\asymp V^{1/4}(\log V)^{\frac{\hat{\theta}}{2}+\hat{\nu}}=V^{1/4}(\log V)^{1/4}.$ (70)
[…]
Appendix D Renormalisation group
In this appendix, we do not work in a mathematically rigorous manner. Instead, we argue as in [28, 29] to compute the profile $f_{n}$ for the (short- or long-range) $n$-component lattice $|\varphi|^{4}$ model in dimensions $d>d_{\mathrm{c},\alpha}=2\alpha$ under PBC. The hypothesis of universality suggests that the same scaling and profile found in this appendix should apply to other $O(n)$ models above the upper critical dimension, including Ising ($n=1$) and XYXY ($n=2$) models.
The $|\varphi|^{4}$ model has spins $\varphi_{x}\in\mathbb{R}^{n}$ and Hamiltonian
$H=\frac{1}{2}(\varphi,(-\Delta)^{\alpha/2}\varphi)+\sum_{x}(\frac{g}{4}|\varphi_{x}|^{4}+\frac{\nu}{2}|\varphi_{x}|^{2}+\vec{h}\cdot\varphi_{x}),$ (72)
where $\vec{h}=(h,\ldots,h)\in\mathbb{R}^{n}$ is a constant external field and $\Delta$ is the discrete Laplacian. We consider both SR (set $\alpha=2$ in all formulas) and LR ($0<\alpha<\min\{2,d\}$) models in dimensions $d>d_{\mathrm{c},\alpha}=2\alpha$. An observable $F$ has expectation
$\langle F\rangle=\frac{1}{Z}\int F(\varphi)e^{-H(\varphi)}D\varphi.$ (73)
The two-point function and susceptibility are
$G_{\nu}(x)=\frac{1}{n}\langle\varphi_{0}\cdot\varphi_{x}\rangle_{h=0},\quad\chi(\nu)=\sum_{x\in\mathbb{Z}^{d}}G_{\nu}(x).$ (74)
For $d\geq 2\alpha$, there is an infinite-volume critical value $\nu_{\mathrm{c}}<0$ such that the Wilson renormalisation group (RG) flow started from $\nu=\nu_{\mathrm{c}}$ converges to a stable Gaussian fixed point.
For PBC, a theory of finite-size scaling based on the RG is developed in [28] and [29, Section 32.3]. That work considered short-range interactions, but it extends mutatis mutandis to long-range, as we present here. We consider now a volume $V=R^{d}$ with PBC in dimensions $d>2\alpha$.
Under the change of scale $x\mapsto L^{-1}x$, the field scales as $\varphi\mapsto L^{-\frac{d-\alpha}{2}}\varphi$. Let $\nu=\nu_{\mathrm{c}}+t$. The model remains near the Gaussian fixed point as long as $L^{\alpha}|t|$ remains bounded. In this regime, briefly put, the free energy is renormalised to linear order as
$f(t,g,h)=L^{-d}f(L^{y_{t}}t,L^{y_{i}}g,L^{y_{h}}h)+\delta f,$ (75)
where $y_{t}=\alpha$, $y_{i}=-(d-2\alpha)$, $y_{h}=\frac{d+\alpha}{2}$, and $\delta f$ is an inhomogeneous term which does not play a role in the present calculation. Near the fixed point, $t(L)\sim a_{d}L^{y_{t}}t$ and $g(L)\sim L^{y_{i}}\bar{g}$ for some constants $a_{d}$ and $\bar{g}$. (We write $f_{1}\sim f_{2}$ for $\lim f_{1}/f_{2}=1$.)
The susceptibility on the torus $\mathbb{T}^{d}_{R}$ can be computed via $\chi_{R}=n^{-1}R^{-d}\frac{\partial^{2}}{\partial h^{2}}\log Z|_{h=0}$. On the torus, the Gaussian field with covariance $((-\Delta)^{\alpha/2})^{-1}$ can be decomposed into two independent Gaussian fields $\zeta+\psi$, where $\psi$ is the constant zero mode and $\zeta$ is the rest of the spin field. Integrating out $\zeta$ corresponds to performing the RG flow to scale $L=R$. Then $\chi_{R}$ is given by the zero-mode integral
$\chi_{R}\sim R^{\alpha}\frac{\int_{\mathbb{R}^{n}}|\psi|^{2}e^{-U_{R}(\psi)}\mathrm{d}\psi}{n\int_{\mathbb{R}^{n}}e^{-U_{R}(\psi)}\mathrm{d}\psi}$ (76)
with
$U_{R}(\psi)=\frac{1}{4}R^{-(d-2\alpha)}\bar{g}|\psi|^{4}+\frac{1}{2}a_{d}R^{\alpha}t|\psi|^{2}.$ (77)
The importance of the zero mode for finite-size scaling under PBC has been stressed in [4] from a different perspective.
After the change of variables $\psi\mapsto\bar{g}^{-1/4}R^{(d-2\alpha)/4}\psi$, and for $\nu=\nu_{\mathrm{c}}+sa_{d}^{-1}\bar{g}^{1/2}R^{-d/2}$ with $s\in(-\infty,\infty)$ (window scale $V^{-1/2}$), this becomes
$\chi_{R}\sim\bar{g}^{-1/2}V^{1/2}f_{n}(s),$ (78)
with the non-Gaussian profile $f_{n}$ given by
$f_{n}(s)=\frac{\int_{\mathbb{R}^{n}}|x|^{2}e^{-\frac{1}{4}x^{4}-\frac{1}{2}sx^{2}}\mathrm{d}x}{\int_{\mathbb{R}^{n}}e^{-\frac{1}{4}x^{4}-\frac{1}{2}sx^{2}}\mathrm{d}x}.$ (79)
After conversion to polar coordinates, this agrees with the formula for $f_{n}$ in (45). With the same value of $\nu$, the torus plateau is given by
$\frac{\chi_{R}}{V}\sim\frac{f_{n}(s)}{\bar{g}^{1/2}V^{1/2}},$ (80)
which dominates the Gaussian contribution $|x|^{-(d-\alpha)}$ to $G_{R}(x)$ when $|x|\gtrsim R^{\frac{d}{2(d-\alpha)}}$; note that the exponent can be written as $(2-d_{\mathrm{c},\alpha}/d)^{-1}$ so is less than $1$. If instead $t=\nu-\nu_{\mathrm{c}}\gg R^{-d/2}$ then a crossover occurs to Gaussian behaviour when $t$ is of order $R^{-\alpha}$. In particular, for such $t$ the susceptibility scales as $R^{\alpha}$, not as $R^{d/2}$. The crossover is discussed in detail for the hierarchical model in [30].
The above identifies the critical window with scale $V^{-1/2}$, the susceptibility scale $V^{1/2}$, the plateau scale $V^{-1/2}$, and the universal profile $f_{n}(s)$, as in Tables 1 and 2.
Universal ratios can be computed similarly. E.g., with the same $\nu=\nu_{\mathrm{c}}+sa_{d}^{-1}\bar{g}^{1/2}V^{-1/2}$, the moments of the total spin $\Phi_{R}=\sum_{x\in\mathbb{T}_{R}^{d}}\varphi_{x}$ obey, as $R\to\infty$,
$\frac{\langle|\Phi_{R}|^{4}\rangle}{\langle|\Phi_{R}|^{2}\rangle^{2}}\to\frac{\int_{\mathbb{R}^{n}}|x|^{4}e^{-\frac{1}{4}x^{4}-\frac{1}{2}sx^{2}}\mathrm{d}x}{(\int_{\mathbb{R}^{n}}|x|^{2}e^{-\frac{1}{4}x^{4}-\frac{1}{2}sx^{2}}\mathrm{d}x)^{2}}.$ (81)
The dimensionless ratio (81) appears in the Binder cumulant and the renormalised coupling constant. At $\nu_{\mathrm{c}}$ ($s=0$), the right-hand side of (81) takes the value
$\frac{n}{4}\left(\frac{\Gamma(\frac{n}{4})}{\Gamma(\frac{n+2}{4})}\right)^{2}.$ (82)